Constants -Part A nter of the as ou If you weigh 690 N on the earth, what would be your weight on the surface of a neutron star that has the same mass as our sun and a diameter of 24.0 km? Take the mass of the sun to be m 1.99-10 kg, the gravitational constant to be G acceleration due to gravity at the earths surface to be g- 9810 m/s 6.67-101N m2/kg2 , and the Express your weight tuitar in newtons View Available Hint(s) Submit Provide Feedback Next>
Answer #1

Weight on the surface of neutron star, W(star)

W(star) = m*g(star)

g(star) = GM/r^2 = 6.67*10^-11*1.99*10^30/(24*10^3/2)^2

= 9.217*10^11

m = 690/9.810 = 70.336 kg

W(star) =70.336*9.217*10^11

= 6.48*10^13 N